# Scalable Dynamic Embedding Size Search for Streaming Recommendation

### v1

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, Idaho, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, Idaho, USAccs: Information systems Recommender systems
@@ 1. Introduction
Many researchers have delved into mixed-dimension embedding size strategies aimed at minimizing memory usage while preserving the performance of recommender systems (Qu et al., 2023; Lyu et al., 2022; Qu et al., 2022; Xia et al., 2022a, 2023a; Liang et al., 2023; Chen et al., 2021; Qu et al., 2024; Liu et al., 2021; Joglekar et al., 2020). These strategies involve assigning different embedding sizes based on the characteristics of users and items, such as their interaction counts. However, these methods are tailored for static and non-streaming settings and the embedding sizes remain fixed after allotment, which largely limits their practical utility in streaming settings for two reasons. Firstly, in streaming environments, where users and items exhibit high dynamism, the embedding sizes need to be continually adjusted to accommodate the evolving user and item behavior. For example, the embedding size of a newly released movie may need to be increased as its frequency rises over time. Maintaining consistent embedding sizes in such a situation will lead to performance degradation because it fails to adopt to the changing frequency levels. Secondly, these methods cannot to assign embedding sizes to newly joined users or items. As a result, implementing these methods in streaming scenarios necessitates retraining them from scratch at the onset of each time segment, introducing significant computational overheads.
@@
  • We conduct comprehensive empirical evaluations on two public datasets, combining various embedding sparsification algorithms with diverse base recommender systems, affirming the advantageous efficacy of SCALL.
@@ 2.3. Embedding Compression in Recommendation
A variety of research endeavors have investigated methods for learning compressed embeddings, serving as a memory-efficient alternative to conventional embedding tables in recommender systems, as outlined in (Zheng et al., 2022; Yin et al., 2024; Tran et al., 2024). Approaches like feature selection search, as exemplified by (Wang et al., 2022; Luo et al., 2019; Lin et al., 2022), focus on training models to learn importance scores for effective filtering of unimportant feature fields. Embedding dimension search methods, illustrated by (Liu et al., 2021; Qu et al., 2022; Kong et al., 2023; Qu et al., 2023, 2024), enhance traditional recommender systems by introducing mixed dimensions for features, departing from uniform sizes. Compositional embedding methods (Liang et al., 2024, 2023; Xia et al., 2023a, b) represents users and items with a smaller number of embedding vectors (aka. meta-embeddings) to shrink the embedding size. Elastic embedding (Chen et al., 2021; Zheng et al., 2024) enable users and items to share embedding blocks. The knowledge distillation method (Xia et al., 2022b) compresses a model by having a smaller student model learn from a larger teacher model. SCALL maps each user and item to their optimal embedding size so it belong to category of embedding dimension search methods.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679638isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
@@ 1. Introduction
Many researchers have delved into mixed-dimension embedding size strategies aimed at minimizing memory usage while preserving the performance of recommender systems (Qu et al., 2023; Lyu et al., 2022; Qu et al., 2022; Liang et al., 2023; Chen et al., 2021; Qu et al., 2024; Liu et al., 2021; Joglekar et al., 2020). These strategies involve assigning different embedding sizes based on the characteristics of users and items, such as their interaction counts. However, these methods are tailored for static and non-streaming settings and the embedding sizes remain fixed after allotment, which largely limits their practical utility in streaming settings for two reasons. Firstly, in streaming environments, where users and items exhibit high dynamism, the embedding sizes need to be continually adjusted to accommodate the evolving user and item behavior. For example, the embedding size of a newly released movie may need to be increased as its frequency rises over time. Maintaining consistent embedding sizes in such a situation will lead to performance degradation because it fails to adopt to the changing frequency levels. Secondly, these methods cannot to assign embedding sizes to newly joined users or items. As a result, implementing these methods in streaming scenarios necessitates retraining them from scratch at the onset of each time segment, introducing significant computational overheads.
@@
  • We conduct comprehensive empirical evaluations on two public datasets and compare SCALL to various embedding sparsification algorithms, affirming the advantageous efficacy of SCALL.
@@ 2.3. Embedding Compression in Recommendation
A variety of research endeavors have investigated methods for learning compressed embeddings, serving as a memory-efficient alternative to conventional embedding tables in recommender systems, as outlined in (Zheng et al., 2022; Yin et al., 2024; Tran et al., 2024). Approaches like feature selection search, as exemplified by (Wang et al., 2022; Luo et al., 2019; Lin et al., 2022), focus on training models to learn importance scores for effective filtering of unimportant feature fields. Embedding dimension search methods, illustrated by (Liu et al., 2021; Qu et al., 2022; Kong et al., 2023; Qu et al., 2023, 2024), enhance traditional recommender systems by introducing mixed dimensions for features, departing from uniform sizes. Compositional embedding methods (Liang et al., 2024, 2023) represents users and items with a smaller number of embedding vectors (aka. meta-embeddings) to shrink the embedding size. Elastic embedding (Chen et al., 2021; Zheng et al., 2024) enable users and items to share embedding blocks. The knowledge distillation method (Xia et al., 2022, 2023) compresses a model by having a smaller student model learn from a larger teacher model. SCALL maps each user and item to their optimal embedding size so it belong to category of embedding dimension search methods.